# Equivalence of scalar-tensor theories and scale-dependent gravity

### v1

@@ V Conclusions
In order to tackle striking issues in modern cosmology like dark matter or dark energy, modified theories of gravity are frequently investigated. While STTs and $f(R)$-gravity likely are the most well-known approaches, there also exist other ideas like SD gravity, which is a low-energy effective theory inspired by asymptotic safety. Studying how such modified gravity theories are related to each other may tell us more about the nature of gravity itself. Therefore, in this article, we have successfully established an equivalence between SD gravity and STTs with a single scalar field, a canonical kinetic term and conformal coupling to the metric tensor. We found that, on the level of the action, any SD theory with smooth running couplings, $G(k)>0$, as well as $B(k)>-\frac{3}{4}\frac{1}{G(k)^{3}}\left(\frac{\partial G(k)}{\partial k}\right)^{2}$ can be uniquely rewritten as a STT. This particularly includes all well-behaved scale-dependent models discussed in the literature, see Ref. Rincón (2019) and references therein, proving that this class of theories can be fully embedded into STTs. Regarding the other direction, there are multiple ways of mapping a STT to an equivalent SD theory, differing in the definitions of the scale-setting relation and the prefactor of the kinetic term for $k$. In this case, the equivalence is applicable to practically any STT whose action is of the form given in Eq. (5). On the level of the field equations, we found that there is an intimate connection between the EOMs for $\phi$ and $k$. In a SD theory resulting from a STT, any $k(\phi)$ will automatically satisfy its EOM. On the other hand, if we start with a SD model in which $k$ does not satisfy its EOM, the scalar field of the equivalent STT will also not satisfy its EOM and must, therefore, be treated as a non-dynamical entity, which is, however, not in the spirit of STTs. This connection between the EOMs lends further support to the idea of promoting $k$ to a dynamical entity with its own kinetic term and EOM. It is important to note that the EOM for $k$ also results from demanding that, in accordance with the Bianchi identities, the divergence of the effective energy-momentum tensor vanishes. In Ref. Reuter and Weyer (2004a), it was shown that by suitably choosing the running couplings and complementing the action with a term containing derivatives of the Newton coupling, a vanishing divergence of the effective energy-momentum tensor can be achieved without imposing the EOM for $k$. However, as we have demonstrated in this article, this additional term in the action is nothing but a kinetic term for $k$ with a specific choice of $B(k)$. By simply treating $k$ as a dynamical field and imposing its EOM, the same result can be achieved for arbitrary running couplings and an arbitrary kinetic term, including $B(k)=0$.
@@
There are some interesting directions that can now be explored. In this work, we have focused mainly on analyzing the equivalence itself, while we have only touched on its applications. Now that the equivalence triangle is completed, any method or result from one theory may be applied to the other two, potentially leading to various new insights about modified gravity in general. Especially, we hope that the relation between the dynamics of the scale-setting relation and the scalar field depicted here motivates the study of SD models, in which $k$ is treated as a dynamical entity. A starting point might be the generalization of the ELFP to the full EOM for $k$ and the subsequent application of this condition to different scenarios, such as cosmology and black holes. Another promising avenue might be to incorporate the equivalence between STTs and certain Higgs portal theories described in Ref. Burrage et al. (2018) into the equivalence picture; see Fig. 2. Since there exist equivalences of STTs with all the aforementioned theories, it can be anticipated that $f(R)$-gravity and SD gravity should also be equivalent to Higgs portal models. Last, there is a crucial aspect of the equivalence that needs further investigation. Even though the action in Eq. (7) is only the lowest order part of the full SD action, it is well-known that this action is a full quantum result and our ignorance of the concrete UV-completion of gravity is incorporated into the running of the gravitational couplings. In sharp contrast, STTs are classical field theories, ignorant of the quantum nature of gravity Laporte et al. (2021). This now raises the question of how an effective quantum gravity theory can be equivalent to a classical field theory, at least at the lowest level of perturbation. Answering this question could lead to new insights into the quantum nature of gravity and STTs.
Figure 2: The equivalence triangle from Fig. 1 amended by equivalences to Higgs portal models Schabinger and Wells (2005); Patt and Wilczek (2006); Ref. Burrage et al. (2018) suggests that every STT of the type considered in the present article can be expressed as a Higgs portal theory, while not every Higgs portal theory is expected to be equivalent to a STT. From this figure, we can conclude that every SD theory with smooth couplings $G(k)>0$ and $B(k)$ fulfilling Eq. (19) should be equivalent to Higgs portal models. However, the direct relations between Higgs portals and SD gravity or $f(R)$-gravity remain to be studied.

### v2

@@ V Conclusions
In order to tackle striking issues in modern cosmology like dark matter or dark energy, modified theories of gravity are frequently investigated. While STTs and $f(R)$-gravity likely are the most well-known approaches, there also exist other ideas like SD gravity, which is a low-energy effective theory inspired by asymptotic safety. Studying how such modified gravity theories are related to each other may tell us more about the nature of gravity itself. Therefore, in this article, we have successfully established an equivalence between SD gravity and STTs with a single scalar field, a canonical kinetic term and conformal coupling to the metric tensor. We found that, on the level of the action, any SD theory with smooth running couplings, $G(k)>0$, as well as $B(k)>-\frac{3}{4}\frac{1}{G(k)^{3}}\left(\frac{\partial G(k)}{\partial k}\right)^{2}$ can be uniquely rewritten as a STT. This particularly includes all well-behaved scale-dependent models discussed in the literature, see Ref. [123] and references therein, proving that this class of theories can be fully embedded into STTs. Regarding the other direction, there are multiple ways of mapping a STT to an equivalent SD theory, differing in the definitions of the scale-setting relation and the prefactor of the kinetic term for $k$. In this case, the equivalence is applicable to practically any STT whose action is of the form given in Eq. (5). On the level of the field equations, we found that there is an intimate connection between the EOMs for $\phi$ and $k$. In a SD theory resulting from a STT, any $k(\phi)$ will automatically satisfy its EOM. On the other hand, if we start with a SD model in which $k$ does not satisfy its EOM, the scalar field of the equivalent STT will also not satisfy its EOM and must, therefore, be treated as a non-dynamical entity, which is, however, not in the spirit of STTs. This connection between the EOMs lends further support to the idea of promoting $k$ to a dynamical entity with its own kinetic term and EOM. It is important to note that the EOM for $k$ also results from demanding that, in accordance with the Bianchi identities, the divergence of the effective energy-momentum tensor vanishes. In Ref. [99], it was shown that by suitably choosing the running couplings and complementing the action with a term containing derivatives of the Newton coupling, a vanishing divergence of the effective energy-momentum tensor can be achieved without imposing the EOM for $k$. However, as we have demonstrated in this article, this additional term in the action is nothing but a kinetic term for $k$ with a specific choice of $B(k)$. By simply treating $k$ as a dynamical field and imposing its EOM, the same result can be achieved for arbitrary running couplings and an arbitrary kinetic term, including $B(k)=0$.
@@
There are some interesting directions that can now be explored. In this work, we have focused mainly on analyzing the equivalence itself, while we have only touched on its applications. Now that the equivalence triangle is completed, any method or result from one theory may be applied to the other two, potentially leading to various new insights about modified gravity in general. Especially, we hope that the relation between the dynamics of the scale-setting relation and the scalar field depicted here motivates the study of SD models, in which $k$ is treated as a dynamical entity. A starting point might be the generalization of the ELFP to the full EOM for $k$ and the subsequent application of this condition to different scenarios, such as cosmology and black holes. Another promising avenue might be to incorporate the equivalence between STTs and certain Higgs portal theories described in Ref. [137] into the equivalence picture; see Fig. 2. Since there exist equivalences of STTs with all the aforementioned theories, it can be anticipated that $f(R)$-gravity and SD gravity should also be equivalent to Higgs portal models. Note that there also exist other investigations of scalar-tensor theories in different conformal frames and parameterizations [138], which have addressed the GR limit [139], inflation [140] and a generalization to the Palatini formalism [141, 142]. It would be interesting to discuss how these studies can be connected to the results obtained in the present article. Last, there is a crucial aspect of the equivalence that needs further investigation. Even though the action in Eq. (7) is only the lowest order part of the full SD action, it is well-known that this action is a full quantum result and our ignorance of the concrete UV-completion of gravity is incorporated into the running of the gravitational couplings. In sharp contrast, STTs are classical field theories, ignorant of the quantum nature of gravity [143]. This now raises the question of how an effective quantum gravity theory can be equivalent to a classical field theory, at least at the lowest level of perturbation. Answering this question could lead to new insights into the quantum nature of gravity and STTs.
Figure 2: The equivalence triangle from Fig. 1 amended by equivalences to Higgs portal models [144, 145]; Ref. [137] suggests that every STT of the type considered in the present article can be expressed as a Higgs portal theory, while not every Higgs portal theory is expected to be equivalent to a STT. From this figure, we can conclude that every SD theory with smooth couplings $G(k)>0$ and $B(k)$ fulfilling Eq. (19) should be equivalent to Higgs portal models. However, the direct relations between Higgs portals and SD gravity or $f(R)$-gravity remain to be studied.